Shift report 20000803 16:00-24:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Thu Aug 03 2000 - 00:01:31 EDT

  • Next message: http account: "Shift report 20000803 00:00-08:00"

    ------------------------------------------------------------
    Date: Thu Aug  3 00:01:31 2000, Shift: 16:00-24:00
    Supervisor: JHL, with: OH
    Log book#3 pages 130 - 132
    
    Shift Summary:
    ------------------------------------------------------------
    1. No useful beam: Working on injection 
    
    2. Magnet interlock (D1, D5) problems fixed by 
       Mr. watch
    
       DB had a quick controlled access, and recycled power of
       FFS1440 (for H1 and C1) and BBL HV: problem fixed.
       And started Isobutane in T4  
    
       Tile HV (MUMF5) Failed: might need recycling power
       (K.O. suspected it's due to humidity?)
     
    3. Nothing new.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Thu Aug 03 2000 - 00:08:06 EDT